An important application of regression analysis in accounting is in the estimation of cost. By collecting data on volume and cost and using the least squares method to develop an estimated regression equation relating volume and cost, an accountant can estimate the cost associated with a particular manufacturing volume. Consider the following sample of production volumes and total cost data for a manufacturing operation.

Production Volume (units) Total Cost ($)
400 4,900
450 5,900
550 6,300
600 6,800
700 7,300
750 7,900


a. Compute b0 and b1 (to 1 decimal).



Complete the estimated regression equation (to 1 decimal).
    

b. What is the variable cost per unit produced (to 1 decimal)?


c. Compute the coefficient of determination (to 3 decimals). Note: report r^2 between 0 and 1.


What percentage of the variation in total cost can be explained by the production volume (to 1 decimal)?


d. The company's production schedule shows 500 units must be produced next month. What is the estimated total cost for this operation (to the nearest whole number)?
Answer #1
Production Volume (X)   Total Cost (Y) X * Y X2 Y2
400 4900 1960000 160000 24010000
450 5900 2655000 202500 34810000
550 6300 3465000 302500 39690000
600 6800 4080000 360000 46240000
700 7300 5110000 490000 53290000
750 7900 5925000 562500 62410000
Total 3450 39100 23195000 2077500 2.6E+08

Part a)

Equation of regression line is Ŷ = a + bX

b = ( 6 * 23195000 - 3450 * 39100 ) / ( 6 * 2077500 - ( 3450 )2)
b = 7.6
a =( Σ Y - ( b * Σ X) ) / n
a =( 39100 - ( 7.6 * 3450 ) ) / 6
a = 2146.7
Equation of regression line becomes Ŷ = 2146.7 + 7.6 X

b0 =  a = 2146.7

b1 = b = 7.6

Part b)

Ŷ = 2146.7 + 7.6 X

Variable cost per unit production is $2146.7.

Part c)



r = 0.979

Coefficient of Determination
R2 = r2 = 0.959

Explained variation = 0.959* 100 = 95.9%

Part d)

When X = 500
Ŷ = 2146.667 + 7.6 X
Ŷ = 2146.667 + ( 7.6 * 500 )
Ŷ = 5946.67 ≈ 5947